Thursday, 23 April 2015

Maintenance Update (lap, face, shiny, fixed)

Friday 17th
Clive reports:
" Regarding the 'hot' pony, neither my driver, Sean Nielsen, or myself noticed any excess heat in that area apart from normal running. Reckon **** has been doing too much washing up without his Marigolds."
{So, I'll write that issue off as investigated + no problem found}

Bruce reports about the clack problem:
"Something wrong here, on Wed we went to quite a lot of trouble to clean the faces and align them properly when clamping them up. What are we doing wrong?"

My theory.
"Let's assume that if both faces are flat, so that when they are clamped up, a THIN gasket should seal ok.  But, as the clamping bolts are outside the raised sealing area it is easy to clamp the faces slightly out of parallel, so, if a thin gasket is used it may not be able to cope with this variation.  Would it be better to use a thicker reinforced gasket to take up any misalignment better?  I acknowledge that a thin gasket is best if you can get it to seal.  Is this valid or am I completely barking up the wrong tree?"

Saturday 18th
David, Bruce and Gilbert spent the whole day working on the leaky clack box joint. Once the clack box was off, Bruce set to, lapping the face to make sure it was perfectly flat.  He found some thicker gasket material and cut a new gasket for it.

Meanwhile, Carpo had looked at the studs that held the clack box to the valve assembly, and decided that the studs were life-expired.  So, David & Gil began removing them.  The usual way is to fit two nuts on the end, tighten them really hard and then undo the pair - usually they grip the threads hard enough to be able to turn the stud and remove it.  This worked on four of the five studs!  After struggling for ages, David decided to weld a nut on the end of the stud and try to turn it that way.  The first attempt ended in the weld breaking before the stud would turn!  So David applied even more weld, and then  finally managed to turn the stud and get it out.

Thereafter, the three of them took it in turns to lap that face.  It's a strain on the arms, because you are moving the lapping block in a vertical plane all of the time.

Carpo made a set of new studs for us {Thank you, Ian!}.  I found some fresh nuts to fit, plus new spring washers.

In a spare moment, Bruce thought to check the nuts on the new blow-down valve, to ensure they were tight after the first few steamings.  They were.

More in hope than expectation, I cleaned the firebox grate and laid a warming fire … but didn't light it!

By end of day, the decision was not to rush things, but complete the job carefully on Wednesday.

Monday 20th
Bruce reported
"This morning David brought his large file to remove material from the Clack Box sealing face to improve the flatness.  We both spent the next couple of hours removing material from the face, but not enough to clean it up.  David then had to leave.

I went back this afternoon and carried on filing, hoping to clean it up, but the flange is so large that I could not remove enough to get rid of the low area completely, but it was getting close.

I then tried lapping the face to get rid of the filing marks and flatten it off and give a better idea how much more needed to be removed.

After several lapping cycles the face had cleaned up except for the low area at the bottom which had only cleaned up across about half its width.

I measured the gap on the bit that had not cleaned up with feeler gauges and found that the largest gap was only 0.003"

In consultation with Carpo I explained that it would probably take at least another half day to clean up the last bit of face completely, so he agreed that it was good enough to reassemble it and see what happens, particularly as we need to be available for the wartime event at the weekend.

It was 5.15pm by the time I had finished so I did not clean the flange.  

Before inserting the studs the threaded holes need cleaning out."

Wednesday 22nd
Jamie & Chris B7 [Loco Dept] took the leaky J-cock off the side of the smokebox and hence found out how it fits!  They gave it a good clean; obtained a new copper washer (from Dinmore Manor group) and reassembled it.  Jamie confirmed that there was oil in the pipe, so that means it is working!

Bruce & Gilbert spent nearly all day working on the clack valve.  The new studs were fitted, but then they didn't quite fit the holes in the clack!  So, Bruce had to tease a couple of holes open a tad.  When it did fit, they verified that nothing was fouling the clack valve and that the two surfaces were absolutely flat against each other.  Bruce also ground a little bit off the clack body to make sure that the nuts & spring washers fitted flush.  The new gasket is thicker and also reinforced.  Carpo kept an eye on things and gave the "all clear" to light a warming fire, which John G and I then did.

Apart from this, Bruce checked the balance weights in the wheels.  John polished everything brass and copper in the cab (to impress the folks on the Fire & Drive on Friday!).  Chris and I lifted the brass bonnet onto the firebox top in anticipation of fitting it after steam test on Thursday.

I had cleaned up four rail chairs during the day, and John painted their bottoms.

Fred was assisted by Geof and Bill at Winchcombe, working on the siphon restoration.

Thursday 23rd
I got to Todders at 10 am.  There were still embers from last night's warming fire, so a quick rake round of the ash, plus a few shovels of coal and I got the fire going again.  However, there was no pressure, and with only half of the grate covered you can't use the blower, so it was after lunch by the time I'd got 30 psi.  Carpo decided to test the injector at that, and there was no sign of a leak.  Then he told me to spread the fire; make a full-grate fire and raise pressure.  By 3 o'clock I'd got 200 psi.  Carpo observed the clack as I injected, and he declared it "passed".

Having got her hot, but going nowhere, she continued to blow off until 4.30 despite periodic injections to cool her down a bit.  Anyway, I fitted the cover over the delivery pipe (to the clack) and then fitted the brass bonnet over the valves.  It was bloomin' hot up there - sun as well as boiler!  I thought my overalls were melting at one point!

I left at 5pm with 180 on the clock; had a take-away; returned at 7pm at which point she had 140 psi.  Checked round; put the chimney cap on; shut the gauge frame; shut the fire hole doors and finally … I'm home!


Thursday, 16 April 2015

Maintenance Update (failed, fixed, failed)

Friday 10th
I knew it was going to happen, but I still have a twinge of sadness - Maurice has definitely left home.  The food I put down for him on Wednesday was still in his little bowl.

2807 was in service on a Fire & Drive.  Seemed to be going OK.  Carpo also told me that he had decided to keep her in service for Saturday.

I really went down to restock the railway shop with some of our teddy bears (the Winchcombe Wendy one seems to be everyone's favourite!).  However, while there I applied Deproma to the tops of the three black-bottomed-chairs.  Then there will be nine ready for their enamel coat on Saturday.

Saturday 11th
Daphne gave Bruce permission to come down for a couple of hours, so he helped me by applying the top coat to the rail chairs.

2807 was chuffing up & down happily, so there was not a lot that people could do - hence I'd suggested they go to Winchcombe and help on the siphon work.

Meanwhile, I tackled some exceedingly clagged-up chairs, and managed only to de-grease four by 2.30 pm.  Bruce had gone by then.  Mike W came and ate lunch in the TPO, but then went to help with P&O (they hope to light a fire on Thursday).

Sunday 12th
2807 was in service again.  I did a spot of painting for a short while, but really went down to count up how many teddy bears we still have.

Tuesday 14th
Carpo reported: " Bad news Roger I'm afraid. The joint twixt safety valve and delivery clack has lasted all of about 4 steamings and has failed again today although the crew managed to nerdle the engine through the day."

Wednesday 15th
Gil & Bruce, assisted by Jeff & John P from Loco Dept, set about solving the problem of why the first gasket lasted four years and the second one only four days.  As the loco was in steam yesterday, it was a shade too warm to work on it initially (not to mention the pressure on the 'clock').

In due course, they removed the clack box from the left-hand side of the safety valve block and retired to the TPO for an inspection.  Bruce started cleaning up the face and it was soon clear that it was not exactly flat!  It was slightly dished, in fact, plus a bit of a dent and a scratch.  So, he and John P spent a while with the diamond scrapey-thing making the surface smooth and flat.

John G & I decided to look at Issue 14, that reported the LH pony running warm.  The loco was not over a pit, so I grovelled I the dirt and John passed tools to & fro.  Draining a small amount of liquid from both axleboxes revealed no water in there (which, of course, there shouldn't be!) only oil.  So we have no idea why it should run warm.  Maybe it was the sunshine?

For most of the rest of the day, John G and I repaired one of the trolleys.  They get (ab)used a lot, and the wooden planks were in a sorry state.  In fact, one piece of plywood had stood up to wear better than real wood.  By chance, there was a heap of plywood planking by the wood store - which is no good for fire lighting.  So we made good use of it!

Once all of the clack box faces were clean and smart, it was refitted by Bruce, Gil & Jeff.  Carpo inspected it and declared Thursday as steam-test day.  Jamie (Loco Dept) is in on Thursday morning and will light a steam-raising fire, so all we needed was someone to light a warming fire tonight ….

Bruce & I had a quick look at the leaky J-cock on the RH side of the smokebox.  Something is loose, but it is hard to be sure exactly what needs doing … and it was still a bit warm inside the smokebox! Anyway, time was running out.

… John assisted me in starting a fire.  By 4.40 pm we decided to adjourn, and I would pop back about 6.30 and bank up the fire.  It is important to only light a small fire when the loco is already hot, otherwise it could easily raise steam during the night and blow-off.  Apart from annoying any neighbours, it could then cause the water level in the boiler to drop to a dangerously low level.

Thursday 16th
FAILED AGAIN!  Water pouring out of the clack box joint.  Yesterday's fix clearly didn't.

Service days
2807 was rostered to be in service every Tuesday, Friday, Saturday and Sunday until the end of the month.  However, this leak will wipe out this weekend.  Otherwise, Wednesdays are good for fixing things!

Speaking with Andrew M (P&O), their planned steaming test for Thursday has been postponed to


Wednesday, 8 April 2015

Maintenance Update (Panic, Don't Panic, clack, steam)

Saturday 4th April
Whilst doing the shopping, Claudette Oddy casually mentioned that "2807 won't be in service today because it failed on Friday".  Panic!!!

Gilbert had gone to Winchcombe to work on the siphon, but toddled down to the platform to watch 2807 come in …. and it didn't!  Panic!!!

Gil shot over to Todders to find out what was happening.  He discovered that she had failed.  He reports:

"2807 failed on Friday after completing two runs to Cheltenham Racecourse.

The failure was caused by a faulty gasket between the LH Feedwater Clack and the Safety Valve base.

The clack valve assembly was removed, the joint faces were cleaned, and a new gasket was cut. During removal of the Clack, one of the fixing studs broke, a new stud was made and inserted into the Valve base. It was noted that at least one of the other studs shows signs of corrosion and will ultimately require replacement. 

During removal of the clack the cover over the top feed pipe was removed and it was noted that the cover is eroding the copper feed pipe. The edge of the cover was dressed to remove any sharp edges in this area. The wear on the pipe will require monitoring.

The clack valve assembly was remounted onto the Safety Valve base and tightened into place. The top feed pipe was reconnected after cleaning the spherical copper washer in the joint. The retaining plate was bolted into place and tightened up.

It is intended to carry out a steam test tomorrow, Sunday, with a view to going back into service.

Thanks are due to "Carpo" and the Loco Dept team, for their work in dealing promptly with this failure, and getting the loco ready for service."

Sunday 5th
Geof popped down to Todders for a short while, and reported:

" The steam test was going well when I left.  Carpo was asking if  any of us could be around tomorrow to put the loco ‘back together’.  I’m available tomorrow from 10.00 am.  Anyone else available?"

on Monday 6th he added:

"Carpo does want to put in  service on Wednesday. Today was cooling down day; tomorrow first thing he will tighten the nuts; then if we put back the water feed pipe covers and dome etc. tomorrow, he will roster it for Wednesday. Hence my request for help; I doubt I should be trying to do this unaided!"

No, Geof, you shouldn't!  [He who had a major heart operation at Christmas!]  Of course, it being Easter, everyone else was well busy with their own things!

Wednesday 8th
2807 was back in service!  Gilbert had managed to attend on Tuesday, and a couple of chaps from Loco Dept did most of the work for us, apparently.  Many thanks to whoever it was [Carpo - please pass on our gratitude].

From what Gil has said, it seems that this failure was something that the boiler inspector had mentioned - that there was a steam leak from the LHS of the safety valve assembly, and it ought to be fixed sooner rather than later  … it has been, now!  Anyway, it is now down as Issue 13:
13 [BE] LHS injector clack leaking (possible joint). Steam seen when injector in use.  { New joint fitted}.

2807 is currently rostered for Friday 10th; Friday 17th to Sunday 19th;  Tuesday 21st; and Friday 24th to Sunday 26th.

Watch out for 2807 on 25th & 26th when it is Wartime in the Cotswolds!
It's always a brilliant event, and if you can make it, do: you will not be disappointed ! 

Oh, I forgot to mention: Wednesday, Fred & Gil were at Winchcombe (mostly), working on the siphon.  John G and I pressed on with boot scrapers - John was head painter; I was needle-gunner and bottom painter!  Bruce passed through briefly to see if we have a diagram of the leaky J-cocks.  Mrs Bruce {Daphne} has a disc problem, and backs do take an awfully long time to get better. So, he's currently excused 2807 duties.


Wednesday, 1 April 2015

Maintenance Update (eccentric, washer, shiny, guide)

Friday 27th
I popped down to Todders to deliver a heap of {Tesco's} cardboard trays that we use to box-up boot scrapers. Mark Y spotted me and pointed out that there's a loose nut & bolt on a bracket that holds oil pipes going down to the axle box on the LHS driving axle.  He & John P had attempted to tighten it, but couldn't.  Mark was concerned that the nut could be catching on the wheel.

Later in the day they had a minor disaster:
The reversing lever was shaking badly and on the second run Mark detected knocking sounds from the motion.  He stopped the loco and went inside the motion to find out what was occurring.  He found that one of the eccentric rods was loose.  Apparently the lock nuts had loosened because the split-pins were not tight up to the locknut.  This had allowed the main nuts to slacken as well, so that the ends of the rod were bumping against the eccentrics.  The nuts were re-tightened and flogged up on return to shed.

We have been advised to get some spacer washers made to take up the space between the locknut and split-pin and to fit them as soon as possible.

It has also been suggested that we check the clearance in the bottom pivot for the reversing lever together with the bushes in the ends of the actuating rod.

Meanwhile:  Maurice has opened up his door to the outside, so he's probably been off out exploring and testing the weather.  Program on TV about calories said that he should only have three peanuts per day!  I better cut back on the bowl of food, then.

Message from Dave Owen (of Cheltenham) via Stuart: "2807 looks a treat and congratulations to all who put in so much time and effort".  It's nice to get feedback like this.

Message from David Holmes in the GWSR Toddington shop: " I now have the Hornby 2807’s in the shop at Toddington"

Saturday 28th
I was late arriving (due to family commitments) but found that I was the first (and only) one to arrive!  The loco was in service anyway, so kind-of hard to do anything on it.

I painted three bottoms; cut up some cardboard boxes, fed Maurice and retired.  Later, I partook of tea & cake in the Coffee Pot cafĂ© on Winchcombe station, and "happened" to see 2807 roll in.  Much comment from onlookers about the cleanliness & shininess of our loco!

Just in case you've never noticed, there's an excellent footplate guide here: that explains the controls.  The cab layout is very similar to that of 2807.

Monday 30th
John G & I went down to see if anything could be done about the issues raised: (a) the eccentric rods, and (b) the oil pipe brackets.

We checked all of the nuts on the eccentrics, and all are well tight.  The space between the locking nuts and their split pins is actually less than the thickness of a washer!  Nevertheless, washers could be machined down to suit.

We went round all of the brackets that hold the oil feed pipes above the axle boxes.  Six of the eight brackets had at least one loose nut!  The problem is that the bolts are countersunk with slotted heads.  This makes it difficult to tighten them … when the nut has seized on the thread!  Also, as the nuts are behind the wheels, and the balance weights were blocking access on the LHS, Carpo had to shunt the loco to give us access to them.  In the end, we had tightened all except two seized nuts.  It was impossible to hold the slotted head firmly enough to turn their nuts … so they won't be dropping off during the week!

Wednesday 1 April
As the loco is in service, no one turned up at Todders.  So, I applied a primer to the tops of the three rail chairs and went home to plant my spuds.

The current list of (new) issues is:

5. [PG]: Regulator suffering from lack of oil?  J-cocks on smokebox leaking.  {Talking with Jeff L, he agreed that he felt that oil was not getting to the regulator when he was driving, and it got worse through the day}

6. [MY] LH engine back gear ecc rod to ecc strap bolt loose.  {He fixed it; I checked the other nuts & bolts}

7. [MY] Crown oil pipe bracket loose on frames, LH driving wheel. Tip of thread catching rear of wheel.  {Monday, John & I went round them all.  Two remain seized.  Not convinced it's catching the wheel}

8. [MY] Vac pump piston rod rubbing on end bush at full forwards stroke.  OK at full rear,  ?Head worn?  {We have been aware that this is so, and we can't do anything about it.  The rod is positioned at its best}

9. [MY] Safety valves lift at 210 psi.  {When talking with Jeff, he suggested that creating a vacuum might become difficult if pressure drops below about 190, so it would be good to adjust them}

10. [AS] Left hand intermediate underkeep retaining bolts too short x 2.  Lock nuts not fully on threads.

11. [IB] Lubricator feed to RH cyls leaking slightly.  Visible on lubricator when hot.

12. [JC] Left hand injector steam feed leaking steam under floor.  {We have been aware that a sleeve is inadequately brazed, but it is a challenge to remove the entire pipe to re-do; so we haven't, yet!}

Loco is in service
Friday 3 April to Monday 6th, then Friday 10th (probably for the weekend … it's just not on the roster yet).